# UniTracker: Learning Universal Whole-Body Motion Tracker for Humanoid Robots

### v2

@@ 5 Conclusion
In this work, we present UniTracker, a unified and scalable framework for whole-body motion tracking in humanoid robots. Built upon a three-stage training pipeline, our approach begins with a privileged teacher policy that enables high-fidelity motion tracking and effective data curation. We then introduce a CVAE-based student policy that achieves robust deployment under partial observations by modeling motion diversity and implicitly incorporating global context. To further extend the system’s adaptability, we propose a lightweight residual decoder for fast adaptation to highly dynamic or out-of-distribution motions. We validate UniTracker extensively in both simulation and real-world settings using a 29-DoF Unitree G1 humanoid. Our method successfully tracks over 8,100 motion sequences with a single policy, outperforming strong teacher-student baselines in terms of accuracy, generalization, and robustness. The results demonstrate the effectiveness of combining generative modeling with hierarchical policy distillation and residual adaptation for expressive and general-purpose humanoid control.

### v3

@@ V Conclusion
In this work, we present UniTracker, a unified and scalable framework for whole-body motion tracking in humanoid robots. Built upon a three-stage training pipeline, our approach begins with a privileged teacher policy that enables high-fidelity motion tracking and effective data curation. We then introduce a CVAE-based student policy that achieves robust deployment under partial observations by modeling motion diversity and implicitly incorporating global context. To further extend the system’s adaptability, we propose a lightweight residual decoder for fast adaptation to highly dynamic or out-of-distribution motions. We validate UniTracker extensively in both simulation and real-world settings using a 29-DoF Unitree G1 humanoid. Our method successfully tracks over 8,100 motion sequences with a single policy, outperforming strong teacher-student baselines and prior methods in terms of accuracy, generalization, and robustness. The results demonstrate the effectiveness of combining generative modeling with hierarchical policy distillation and residual adaptation for expressive and general-purpose humanoid control.